# Triality and Adjoint lifting for ${\rm GL}_{3}$

### v1

@@ 1. Introduction
Let $\pi$ be a unitary cuspidal automorphic representation of ${\rm GL}_{3}$ and assume that ${\rm Ad}(\pi)$ has been constructed as the composite of the Langlands functorial lifting given in (1.1), as discussed above. Then ${\rm Ad}(\pi)$ is a self-dual automorphic representation of orthogonal type with trivial central character. Moreover, one has:
@@
Acknowledgments: We thank Gaetan Chenevier (who should have been a coauthor of this paper), Erez Lapid, Dipendra Prasad, Sug Woo Shin and Yugo Takanashi for helpful discussions at various stages of this work and comments on earlier drafts of this paper. In addition, we thank Henry Kim for sending us a copy of the paper [KS3] and Freydoon Shahidi for discussion on the fibers of the Rankin-Selberg lifting from ${\rm GL}_{2}\times{\rm GL}_{3}$ to ${\rm GL}_{6}$. The author is supported by a Tan Chin Tuan Centennial Professorship at NUS.
@@ 7.1. Global K-forms of ${\rm PGSO}_{8}$
Let $\pi$ be a cuspidal representation of ${\rm GL}_{3}$ such that $\pi_{v_{0}}$ is a discrete series representation for some place $v_{0}$. Then the (weak) adjoint lifting of $\pi$ to ${\rm GL}_{8}$ exists.
@@
Now we may use the action of the spherical Hecke algebras of $G={\rm PGSO}_{8}$ at almost all places of $k$ to isolate a given Hecke-Satake family on the LHS of (7.4). More precisely, with our given cuspidal representation $\pi$ of ${\rm GL}_{3}$ with associated submodule $L[\pi]\subset\mathcal{A}_{cusp}({\rm SL}_{3})$ and Hecke-Satake family
@@
where $\tilde{\rm Ad}:{\rm PGL}_{3}({\mathbb{C}})\hookrightarrow{\rm Spin}_{8}({\mathbb{C}})$ is the adjoint map and $S$ is a sufficiently large finite set of places, including all Archimedean ones and the place $v_{0}$. At each Archimedean place $v$, we may also consider the infinitesimal character $\lambda([\pi_{v}])$ of $\mathcal{L}[\pi_{v}]$, which may be regarded as a semisimple conjugacy class in the complexified dual Lie algebra $\mathfrak{sl}_{3}$, giving rise to
@@
(7.5) $\sum_{\mathbb{O}}I_{disc,c,\lambda}^{\mathbb{O},\theta}(f_{\mathbb{O},S})=\sum_{[\pi^{\prime}]_{s}}S^{{\rm SL}_{3}}_{disc,[\pi^{\prime}]_{s}}(f_{{\rm SL}_{3},S})$
@@
with $1_{K_{v}}$ the characteristic function of $K_{v}={\rm PGSO}_{8}(\mathcal{O}_{v})$, and the sum on the RHS ranges over all $\pi^{\prime}]_{s}$ such that
@@
We shall denote the set of such $\pi^{\prime}]_{s}$ by $\Xi_{c,\lambda}$.
Now for each $\pi^{\prime}]\in\Xi_{c,\lambda}$, it follows by Proposition 5.8 that, for each $v\notin S$, the local L-packets $\mathcal{L}[\pi^{\prime}_{v}]$ and $\mathcal{L}[\pi_{v}]$ are either equal or contragredient of each other; in particular, $\mathcal{L}[\pi^{\prime}_{v}]$ is a generic L-packet for each $v\notin S$. Moreover, since $f_{v_{0}}$ lies in $\mathcal{I}^{\theta}_{cusp}({\rm PGSO}_{8})$, any $\pi^{\prime}]_{s}\in\Xi_{c,\lambda}$ is necessarily contained in the cuspidal spectrum $\mathcal{A}_{cusp}({\rm SL}_{3})$.
@@
$\sum_{\mathbb{O}}I_{disc,c}^{\mathbb{O},\theta}(f_{\mathbb{O},S})=\sum_{[\pi^{\prime}]_{s}\in\Xi_{c,\lambda}}{\rm Tr}\left(\pi^{\prime}_{S}(f_{{\rm SL}_{3},S})\right).$
@@
$\sum_{\mathbb{O}}I_{disc,c}^{\mathbb{O},\theta}\neq 0.$
This shows that the Hecke-Satake family $c$ can be detected in the spectrum of the $I_{disc}^{\mathbb{O},\theta}$ for some $\mathbb{O}$ and hence can be realized by an automorphic subrepresentation $\Pi_{\mathbb{O}}$ of ${\rm PGSO}_{8}^{\mathbb{O}}$ for some $\mathbb{O}$.
Pulling $\Pi_{\mathbb{O}}$ back to ${\rm SO}_{8}$ via $f_{1}:{\rm SO}_{8}^{\mathbb{O}}\rightarrow{\rm PGSO}_{8}^{\mathbb{O}}$ and transfering of $f_{1}^{*}(\Pi_{\mathbb{O}})$ to ${\rm GL}_{8}$ (appealing to, for example, [CZ] or [CFK] for the case when $\mathbb{O}$ is not split), one obtains the desired weak adjoint lifting of $\pi$. ∎
@@ 8. Automorphic Induction Case
In this section, we study the case when the cuspidal representation $\pi$ is obtained by automorphic induction from a Hecke character of a cubic field extension $E/k$. In particular, we shall see that the adjoint lifting ${\rm Ad}(\pi)$ always exists (without the local condition in Theorem 7.2. Moreover, we shall be able to determine the possible isobaric decomposition of ${\rm Ad}(\pi)$.
@@ 11.2. Heisenberg-Weil representation
By the theory of Weil representations, this representation can be extended to to ${\rm Sp}(W)$. Since ${\rm Sp}(W)={\rm SL}_{2}({\mathbb{Z}}/3{\mathbb{Z}})$ is not perfect, the extension is of course not unique, as we may twist by cubic characters of ${\rm Sp}(W)$. However, the restriction of these extensions to the commutator subgroup $Q_{8}$ are isomorphic. We will denote this extended representation of $H(W)\rtimes Q_{8}$ by $\omega_{\chi}$ as well, and call it the Heisenberg-Weil representation. The restriction of $\omega_{\chi}$ to each subgroup of $H(W)\rtimes Q_{8}$ is thus irreducible.
@@
Let $M/F$ be a Galois extension with Galois group isomorphic to the Heisenberg group $H(W)$ and consider its irreducible 3-dimensional representation $\omega_{\chi}$ constructed above. The fixed field of the center $Z$ of $H(W)$ is thus a bi-cyclic-cubic extension $E_{Z}$ of $F$. For any line $L\subset W$, the fixed field of the abelian normal subgroup $H(L)$ is a Galois cubic extension $E_{L}$ of $F$. Since
@@
$\pi={\rm AI}_{E_{L}/F}(\chi_{L}).$
In other words, $\pi$ is an automorphic induction from $E_{L}$. Since there are 4 possible lines $L\subset W$ (as $|\mathbb{P}^{1}(\mathbb{F}_{3})|=4$), we see that $\pi$ is an automorphic induction from the 4 cubic subfields in the bi-cyclic-cubic extension $E_{Z}$ of $F$, so that we are in situation (b) of Theorem 8.2. In particular, ${\rm Ad}(\pi)$ is the sum of the 8 cubic characters $\omega_{E_{L}/F}^{\pm 1}$.
@@
Now suppose that $M/F$ is a Galois extension with Galois group $H(W)\rtimes\mu_{2}$. Let $K$ be the fixed field of $H(W)$, so that $K/F$ is a quadratic extension. The action of ${\rm Gal}(K/F)$ on Gal(M/K){Gal}(M/K) is the action of $\mu_{2}\subset{\rm Sp}(W)$ on $H(W)$. In particular, it acts as 1-1 on ${\rm Gal}(E_{Z}/K)=W$ (where $E_{Z}$ is the fixed field of the center $Z$ of $H(W)$). This shows that for each line $L\subset W$, its fixed field $E_{L}$ is a cubic extension of $K$, which is Galois over $F$ with ${\rm Gal}(E_{L}/F)\simeq S_{3}$.
Now the irreducible 3-dimensional representation $\omega_{\chi}$ of $H(W)$ gives rise to a cuspidal representation $\pi_{K}$ of ${\rm GL}_{3}({\mathbb{A}}_{K})$ as described above. Hence, $\pi_{K}$ is an automorphic induction from any of the 4 cubic subfields $E_{L}$ over $K$. Since $\omega_{\chi}$ extends to $H(W)\rtimes\mu_{2}$, the cuspidal representation $\pi_{K}$ is invariant under the action of ${\rm Gal}(K/F)$; in particular, it follows from [AC] that $\pi_{K}$ is the base change of a cuspidal representation $\pi$ of ${\rm GL}_{3}$ over $F$. Moreover, since ${\rm Ad}(\pi_{K})$ is the isobaric sum of the 8 cubic characters $\omega_{E_{L}/K}^{\pm 1}$, we deduce that ${\rm Ad}(\pi)$ is the isobaric sum of 4 cuspidal representations of ${\rm GL}_{2}$. In particular, we are in the situation of Theorem 8.4(c).
@@
Finally, suppose that $M/F$ is a Galois extension with Galois group $H(W)\rtimes C_{4}$ and set
@@
Then $L/F$ is a cyclic degree $4$ extension with Galois group $C_{4}$. By the previous subsection, the irreducible representation $\omega_{\chi}$ of $H(W)\rtimes\mu_{2}$ gives rise to a cuspidal representation $\pi_{K}$ of ${\rm GL}_{3}({\mathbb{A}}_{K})$ such that ${\rm Ad}(\pi_{K})$ is the isobaric sum of 4 cuspidal representations of ${\rm GL}_{2}({\mathbb{A}}_{K})$, corresponding to the 4 lines in $W$.
Since $\omega_{\chi}$ extends to $H(W)\rtimes C_{4}$, we see that $\pi_{K}$ is invariant under ${\rm Gal}(K/F)$, and hence is the base change of a cuspidal representation of ${\rm GL}_{3}$ over $F$ by [AC]. Moreover, the action of ${\rm Gal}(L/F)=C_{4}$ on ${\rm Gal}(E_{Z}/L)=W$ (where $E_{Z}$ is the fixed field of the center $Z\subset H(W)$) is generated by the action of the Weyl element
@@
As the element $w_{0}^{2}$ acts as 1-1 on $W$, we see that the action of $C_{4}$ has two orbits on the set of lines in $W$. This shows that ${\rm Gal}(K/F)$ has two orbits on the 4 isobaric summands in ${\rm Ad}(\pi_{K})$, and hence ${\rm Ad}(\pi)$ is the isobaric sum of two cuspidal representations of ${\rm GL}_{4}$ over $F$. In particular, we are in the situation of Theorem 10.1(ii).
@@
Not surprisingly, if we consider a Galois extension with Galois group $H(W)\rtimes Q_{8}$, a similar argument as above will lead to a cuspidal $\pi$ such that ${\rm Ad}(\pi)$ is cuspidal on ${\rm GL}_{8}$ (i.e. the situation of Theorem 10.1(i)) but the base change of ${\rm Ad}(\pi)$ to an appropriate quadratic extension decomposes as an isobaric sum of two cuspidal ${\rm GL}_{4}$-summands (i.e. the situation of Theorem 10.1(ii)).
We note that in closing that, by a result of Lapid [L2, §5, Prop. 2], the existence of the adjoint lifting as shown in Theorem 7.2 implies that the Galois representation $\omega_{\chi}$ of $H(W)\rtimes\mu_{2}$, $H(W)\rtimes C_{4}$ or $H(W)\rtimes Q_{8}$ corresponds to the cuspidal representation $\pi$ in the respective cases under the purported global Langlands correspondence.

### v2

@@ 1. Introduction
Let $\pi$ be a unitary cuspidal automorphic representation of ${\rm GL}_{3}$ and assume that ${\rm Ad}(\pi)$ has been constructed as the composite of the Langlands functorial lifting given in (1.1), as discussed above. Then ${\rm Ad}(\pi)$ is an isobaric sum of cuspidal representations with unitary central characters. Moreover, ${\rm Ad}(\pi)$ is self-dual of orthogonal type with trivial central character. More precisely, one has:
@@
As a final result for this paper, we exploit the existence of the adjoint lifting to show the following bound towards the Ramanujan-Petersson conjecture for ${\rm GL}_{3}$:
Theorem 1.4.
Let $\pi$ be a cuspidal representation of ${\rm GL}_{3}$ with unitary central character and a discrete series local component over a number field $k$. At each place $v$ where $\pi_{v}$ is unramified, let $\lambda$ be an eigenvalue of its Satake parameter in ${\rm GL}_{3}({\mathbb{C}})$. Writing $|\lambda|=q_{v}^{a}$ for some $a\in{\mathbb{R}}$ (and with $q_{v}$ equal to the cardinality of the residue field of $k_{v}$), one has:
$|a|\leq\frac{21}{65}<\frac{1}{3}.$
This relies on the general bounds of Luo-Rudnick-Sarnak [LRS] and improves upon a result of Blomer-Brumley [BB], which gives $|a|\leq 5/14$ for ${\rm GL}_{3}$. The Ramanujan-Petersson conjecture, of course, asserts that $a=0$. A recent paper of LY Yang [Y], extending earlier work of Ramakrishnan [Ra3] over $\mathbb{Q}$, shows that this holds for infinitely many places $v$ of $k$.
Acknowledgments: We thank Gaetan Chenevier (who should have been a coauthor of this paper), Erez Lapid, Dipendra Prasad, Sug Woo Shin and Yugo Takanashi for helpful discussions at various stages of this work and comments on earlier drafts of this paper. In addition, we thank Henry Kim for sending us a copy of the paper [KS3], Freydoon Shahidi for discussion on the fibers of the Rankin-Selberg lifting from ${\rm GL}_{2}\times{\rm GL}_{3}$ to ${\rm GL}_{6}$, and David Hansen for his catalytic question which leads to the content of Section 12 and the proof of Theorem 1.4. The author is supported by a Tan Chin Tuan Centennial Professorship at NUS.
@@ 7.1. Global K-forms of ${\rm PGSO}_{8}$
Let $\pi$ be a cuspidal representation of ${\rm GL}_{3}$ such that $\pi_{v_{0}}$ is a discrete series representation for some place $v_{0}$. Then the (weak) adjoint lifting of $\pi$ exists as an automorphic representation ${\rm Ad}(\pi)$ of ${\rm GL}_{8}$. Moreover, if $\pi_{v}$ is unramified with Satake parameter $c(\pi_{v})\in{\rm GL}_{3}({\mathbb{C}})$, then ${\rm Ad}(\pi)_{v}$ is unramified with Satake parameter ${\rm Ad}(c(\pi_{v}))\in{\rm GL}_{8}({\mathbb{C}})$.
@@
Now we may use the action of the spherical Hecke algebras of $G={\rm PGSO}_{8}$ at almost all places of $k$ to isolate a given Hecke-Satake family on the LHS of (7.4). More precisely, with our given cuspidal representation $\pi$ of ${\rm GL}_{3}$ with associated submodule $L[\pi]\subset\mathcal{A}_{cusp}({\rm SL}_{3})$, let $S$ be the finite set of places (containing all Archimedean places and the place $v_{0}$) such that $\pi_{v}$ is unramified if and only if $v\notin S$. Considering the Hecke-Satake family
@@
where $\tilde{\rm Ad}:{\rm PGL}_{3}({\mathbb{C}})\hookrightarrow{\rm Spin}_{8}({\mathbb{C}})$ is the adjoint map. At each Archimedean place $v$, we may also consider the infinitesimal character $\lambda([\pi_{v}])$ of $\mathcal{L}[\pi_{v}]$, which may be regarded as a semisimple conjugacy class in the complexified dual Lie algebra $\mathfrak{sl}_{3}$, giving rise to
@@
(7.5) $\sum_{\mathbb{O}}I_{disc,S,c,\lambda}^{\mathbb{O},\theta}(f_{\mathbb{O},S})=\sum_{[\pi^{\prime}]_{s}}S^{{\rm SL}_{3}}_{disc,[\pi^{\prime}]_{s}}(f_{{\rm SL}_{3},S})$
@@
with $1_{K_{v}}$ the characteristic function of $K_{v}={\rm PGSO}_{8}(\mathcal{O}_{v})$, and the sum on the RHS ranges over all $\pi^{\prime}]_{s}$ whose restriction to ${\rm SL}_{3}$ is unramified outside $S$, and such that
@@
We shall denote the set of such $\pi^{\prime}]_{s}$ by $\Xi_{S,c,\lambda}$.
Now for each $\pi^{\prime}]\in\Xi_{S,c,\lambda}$, it follows by Proposition 5.8 that, for each $v\notin S$, the local L-packets $\mathcal{L}[\pi^{\prime}_{v}]$ and $\mathcal{L}[\pi_{v}]$ are either equal or contragredient of each other; in particular, $\mathcal{L}[\pi^{\prime}_{v}]$ is a generic L-packet for each $v\notin S$. Moreover, since $f_{v_{0}}$ lies in $\mathcal{I}^{\theta}_{cusp}({\rm PGSO}_{8})$, any $\pi^{\prime}]_{s}\in\Xi_{S,c,\lambda}$ is necessarily contained in the cuspidal spectrum $\mathcal{A}_{cusp}({\rm SL}_{3})$.
@@
$\sum_{\mathbb{O}}I_{disc,S,c,\lambda}^{\mathbb{O},\theta}(f_{\mathbb{O},S})=\sum_{[\pi^{\prime}]_{s}\in\Xi_{S,c,\lambda}}{\rm Tr}\left(\pi^{\prime}_{S}(f_{{\rm SL}_{3},S})\right).$
@@
$\sum_{\mathbb{O}}I_{disc,S,c,\lambda}^{\mathbb{O},\theta}\neq 0.$
This shows that the pair $(c,\lambda)$ can be detected in the spectrum of the $I_{disc}^{\mathbb{O},\theta}$ for some $\mathbb{O}$ and hence can be realized by an automorphic subrepresentation $\Pi_{\mathbb{O}}$ of ${\rm PGSO}_{8}^{\mathbb{O}}$ for some $\mathbb{O}$. Moreover, such a $\Pi_{\mathbb{O}}$ is unramified outside of $S$.
Pulling $\Pi_{\mathbb{O}}$ back to ${\rm SO}_{8}$ via $f_{1}:{\rm SO}_{8}^{\mathbb{O}}\rightarrow{\rm PGSO}_{8}^{\mathbb{O}}$ and transfering of $f_{1}^{*}(\Pi_{\mathbb{O}})$ to ${\rm GL}_{8}$ (appealing to, for example, [CZ] or [CFK] for the case when $\mathbb{O}$ is not split), one obtains the desired weak adjoint lifting ${\rm Ad}(\pi)$ of $\pi$. More precisely, by its construction, ${\rm Ad}(\pi)$ is unramified outside of $S$ with Satake oparameter ${\rm Ad}(c(\pi_{v}))$ for all $v\notin S$, and with infinitesimal character ${\rm Ad}_{*}(\lambda[\pi_{\infty}])\in\mathfrak{gl}_{8}(k\otimes_{\mathbb{Q}}{\mathbb{C}})$. ∎
@@ 8. Automorphic Induction Case
The next few sections of the paper are devoted to the analysis of ${\rm Ad}(\pi)$ and in particular to the proof of Theorem 1.3. In this section, we examine the case when the cuspidal representation $\pi$ is obtained by automorphic induction from a Hecke character of a cubic field extension $E/k$. In particular, we shall see that the adjoint lifting ${\rm Ad}(\pi)$ always exists (without the local condition in Theorem 7.2. Moreover, we shall be able to determine the possible isobaric decomposition of ${\rm Ad}(\pi)$.
@@ 11.2. Heisenberg-Weil representation
By the theory of Weil representations, this representation can be extended to to ${\rm Sp}(W)$. Since ${\rm Sp}(W)={\rm SL}_{2}({\mathbb{Z}}/3{\mathbb{Z}})$ is not perfect, the extension is of course not unique, as we may twist by cubic characters of ${\rm Sp}(W)$. However, the restriction of these extensions to the commutator subgroup $Q_{8}$ are isomorphic. We will denote this extended representation of $H(W)\rtimes Q_{8}$ by $\omega_{\chi}$ as well, and call it the Heisenberg-Weil representation. The restriction of $\omega_{\chi}$ to any subgroup of $H(W)\rtimes Q_{8}$ containing $H(W)$ is thus irreducible.
@@
Let $M/k$ be a Galois extension with Galois group isomorphic to the Heisenberg group $H(W)$ and consider its irreducible 3-dimensional representation $\omega_{\chi}$ constructed above. The fixed field of the center $Z$ of $H(W)$ is thus a bi-cyclic-cubic extension $E_{Z}$ of $k$. For any line $L\subset W$, the fixed field of the abelian normal subgroup $H(L)$ is a Galois cubic extension $E_{L}$ of $k$. Since
@@
$\pi={\rm AI}_{E_{L}/k}(\chi_{L}).$
In other words, $\pi$ is an automorphic induction from $E_{L}$. Since there are 4 possible lines $L\subset W$ (as $|\mathbb{P}^{1}(\mathbb{F}_{3})|=4$), we see that $\pi$ is an automorphic induction from the 4 cubic subfields in the bi-cyclic-cubic extension $E_{Z}$ of $k$, so that we are in situation (b) of Theorem 8.2. In particular, ${\rm Ad}(\pi)$ is the sum of the 8 cubic characters $\omega_{E_{L}/k}^{\pm 1}$.
@@
Now suppose that $M/k$ is a Galois extension with Galois group $H(W)\rtimes\mu_{2}$. Let $K$ be the fixed field of $H(W)$, so that $K/k$ is a quadratic extension. If $E_{Z}$ is the fixed field of the center $Z$ of $H(W)$, then the action of ${\rm Gal}(K/k)$ on ${\rm Gal}(E_{Z}/K)=W$ is the action of $\mu_{2}\subset{\rm Sp}(W)$ on $W$. In other words, it acts as 1-1 on ${\rm Gal}(E_{Z}/K)=W$. This shows that for each line $L\subset W$, its fixed field $E_{L}$ is a cubic extension of $K$, which is Galois over $k$ with ${\rm Gal}(E_{L}/k)\simeq S_{3}$.
Now the irreducible 3-dimensional representation $\omega_{\chi}$ of $H(W)$ gives rise to a cuspidal representation $\pi_{K}$ of ${\rm GL}_{3}({\mathbb{A}}_{K})$ as described above. Hence, $\pi_{K}$ is an automorphic induction from any of the 4 cubic subfields $E_{L}$ of $M/K$. Since $\omega_{\chi}$ extends to $H(W)\rtimes\mu_{2}$, the cuspidal representation $\pi_{K}$ is invariant under the action of ${\rm Gal}(K/k)$; in particular, it follows from [AC] that $\pi_{K}$ is the base change of a cuspidal representation $\pi$ of ${\rm GL}_{3}$ over $k$. Now the cuspidal representation $\pi$ is not unique, as we can twist it by the quadratic Hecke character $\omega_{K/k}$. But there is a unique such $\pi$ whose central character $\omega_{\pi}$ corresponds to $\det\omega_{\chi}$ under global class field theory. By [L2, §5, Prop. 2], this $\pi$ then corresponds to $\omega_{\chi}$ under the purported global Langlands correspondence for ${\rm GL}_{3}$. Moreover, since ${\rm Ad}(\pi_{K})$ is the isobaric sum of the 8 cubic characters $\omega_{E_{L}/K}^{\pm 1}$, we deduce that ${\rm Ad}(\pi)$ is the isobaric sum of 4 cuspidal representations of ${\rm GL}_{2}$. In particular, we are in the situation of Theorem 8.4(c).
@@
Finally, suppose that $M/k$ is a Galois extension with Galois group $H(W)\rtimes C_{4}$ and set
@@
Then $L/k$ is a cyclic degree $4$ extension with Galois group $C_{4}$. By the previous subsection, the irreducible representation $\omega_{\chi}$ of $H(W)\rtimes\mu_{2}$ gives rise to a cuspidal representation $\pi_{K}$ of ${\rm GL}_{3}({\mathbb{A}}_{K})$ such that ${\rm Ad}(\pi_{K})$ is the isobaric sum of 4 cuspidal representations of ${\rm GL}_{2}({\mathbb{A}}_{K})$, corresponding to the 4 lines in $W$.
Since $\omega_{\chi}$ extends to $H(W)\rtimes C_{4}$, we see that $\pi_{K}$ is invariant under ${\rm Gal}(K/F)$, and hence is the base change of a cuspidal representation $\pi$ of ${\rm GL}_{3}$ over $k$ by [AC]. As in the previous subsection, there is a unique such $\pi$ if we require that its central character corresponds to $\det\omega_{\chi}$ under global class field theory. By [L2, §5, Prop. 2] again, one knows that this unique $\pi$ corresponds to $\omega_{\chi}$ under the purported global Langlands correspondence.
Now the action of ${\rm Gal}(L/k)=C_{4}$ on ${\rm Gal}(E_{Z}/L)=W$ (where $E_{Z}$ is the fixed field of the center $Z\subset H(W)$) is generated by the action of the Weyl element
@@
As the element $w_{0}^{2}$ acts as 1-1 on $W$, we see that the action of $C_{4}$ has two orbits on the set of lines in $W$. This shows that ${\rm Gal}(K/k)$ has two orbits on the 4 isobaric summands in ${\rm Ad}(\pi_{K})$, and hence ${\rm Ad}(\pi)$ is the isobaric sum of two cuspidal representations of ${\rm GL}_{4}$ over $k$. In particular, we are in the situation of Theorem 10.1(ii).
@@
Not surprisingly, if we consider a Galois extension with Galois group $H(W)\rtimes Q_{8}$, a similar argument as above will lead to a cuspidal $\pi$ such that ${\rm Ad}(\pi)$ is cuspidal on ${\rm GL}_{8}$ (i.e. the situation of Theorem 10.1(i)) but the base change of ${\rm Ad}(\pi)$ to an appropriate quadratic extension decomposes as an isobaric sum of two cuspidal ${\rm GL}_{4}$-summands. This gives an instance of Theorem 10.1(ii), though one might expect the general situation of Theorem 10.1(ii) will be such that ${\rm Ad}(\pi)$ is not invariant under twisting by any quadratic Hecke character.
12. Application to Ramanujan Bounds
In this final section, we give an application of the existence of the adjoint lifting to the question of bounding the Hecke eigenvalues of cuspidal representations of ${\rm GL}_{3}$.
12.1. Ramanujan-Petersson conjecture
Let $\pi$ be a cuspidal representation of ${\rm GL}_{n}$ over $k$ with unitary central character. A a place $v$ of $k$ where $\pi_{v}$ is unramified, the Satake parameter $c(\pi_{v})$ is a semisimple conjugacy class in ${\rm GL}_{n}({\mathbb{C}})$. For any eigenvalue $\lambda$ of $c(\pi_{v})$, we may write
$|\lambda|=q_{v}^{a}$
where $q_{v}$ is the cardinality of the residue field of $k_{v}$ and $a\in{\mathbb{R}}$. The Ramanujan-Petersson conjecture asserts that $|\lambda|=1$ so that $a=0$.
12.2. Some prior results
A general bound towards the Ramanujan-Petersson conjecture was established by Luo-Rudnick-Sarnak [LRS]. They showed that
$|a|\leq\frac{1}{2}-\frac{1}{n^{2}+1},$
so that for the case of ${\rm GL}_{3}$, this reads:
$|a|\leq\frac{1}{2}-\frac{1}{10}=\frac{2}{5}.$
There have certainly been improvements to this general bound when $n$ is small. In particular, for the case of ${\rm GL}_{3}$, the paper [BB] of Blomer-Brumley showed that
$|a|\leq\frac{5}{14},$
thus improving upon the Luo-Rudnick-Sarnak bound.
12.3. Application of adjoint lifting
As an application of the existence of the adjoint lifting ${\rm Ad}(\pi)$ (as supplied by Theorem 7.2), we shall show:
Theorem 12.1.
Let $\pi$ be a cuspidal representation of ${\rm GL}_{3}$ with unitary central character and a discrete series local component over a number field $k$. At each place $v$ where $\pi_{v}$ is unramified, let $\lambda$ be an eigenvalue of its Satake parameter $c(\pi_{v})\in{\rm GL}_{3}({\mathbb{C}})$, and write $|\lambda|=q_{v}^{a}$ for some $a\in{\mathbb{R}}$. Then
$|a|\leq\frac{21}{65}<\frac{1}{3}.$
Proof.
By Theorem 7.2, ${\rm Ad}(\pi)$ exists as an automorphic representation of ${\rm GL}_{8}$. Moreover, by Theorem 1.3, ${\rm Ad}(\pi)$ is an isobaric sum of cuspidal representations (of ${\rm GL}_{r}$, with $r\leq 8$) with unitary central characters.
Let us fix the unramified place $v$ and let $\lambda_{i}$ (for $1\leq i\leq.3)$ be the three eigenvalues of the Satake parameter $c(\pi_{v})$. Writing $|\lambda_{i}|=p^{a_{i}}$, the condition of unitary central character implies that
$a_{1}+a_{2}+a_{3}=0.$
On the other hand, the Satake parameter of ${\rm Ad}(\pi)_{v}$ has eigenvalues $\lambda_{i}/\lambda_{j}$, with $|\lambda_{i}/\lambda_{j}|=q_{v}^{a_{i}-a_{j}}$. Applying the Luo-Rudnick-Sarnak bound to each isobaric summand of ${\rm Ad}(\pi)$, one has
$|a_{i}-a_{j}|\leq\frac{1}{2}-\frac{1}{65}.$
We would like to exploit this system of inequalities to get the desired bound on $|a_{i}|$.
For this purpose, note that
$3a_{1}=(a_{1}+a_{2}+a_{3})+(a_{1}-a_{2})+(a_{1}-a_{3}).$
The triangle inequality thus implies that
$3\cdot|a_{1}|\leq|a_{1}+a_{2}+a_{3}|+|a_{1}-a_{2}|+|a_{1}-a_{3}|\leq 1-\frac{2}{65},$
and hence
$|a_{1}|\leq\frac{1}{3}-\frac{2}{195}=\frac{21}{65}<\frac{1}{3}.$
The same argument applies to give the same bound for $|a_{2}|$ and $|a_{3}|$. This proves the theorem. ∎
The following corollary of the theorem was mentioned to us by D. Prasad:
Corollary 12.2.
In the context of Theorem 12.1, let $|\lambda_{i}|=q_{v}^{a_{i}}$ (for $1\leq i\leq 3$) be the absolute value of the three eigenvalues of the Satake parameter of $\pi_{v}$. Then
$|a_{1}|+|a_{2}|+|a_{3}|\leq\frac{42}{65}.$
Proof.
Without loss of generality, we may assume that $a_{1}\geq a_{2}\geq a_{3}$. Since $a_{1}+a_{2}+a_{3}=0$, we have one of the following two possibilities:
$a_{1}\geq a_{2}\geq 0\geq a_{3}\quad\text{or}\quad a_{1}\geq 0\geq a_{2}\geq a_{3}.$
In the first scenario, one has:
$|a_{1}|+|a_{2}|+|a_{3}|=a_{1}+a_{2}-a_{3}=-2a_{3}=2|a_{3}|$
whereas in the second,
$|a_{1}|+|a_{2}|+|a_{3}|=a_{1}-a_{2}-a_{3}=2a_{1}=2|a_{1}|.$
The desired inequality thus follows by Theorem 12.1. ∎